OpenStudy (anonymous):

f(x)=2x+4 g(x)=3x^2-2x-3 Find f(-6) and g(3)

OpenStudy (anonymous):

Just sub -6 and 3 to each equation!

OpenStudy (anonymous):

f(x)=2x+4 g(x)=3x^2-2x=3 f(-6)=2(-6)+4 =-12+4 =-8 g(3)=3(3)^2-2(3)-3 =3(9)-6-3 =27-9 =16